# LEGO: Learning and Graph-Optimized Modular Tracker for Online Multi-Object Tracking with Point Clouds

### v4

@@ Abstract
Online multi-object tracking (MOT) plays a pivotal role in autonomous systems. The state-of-the-art approaches usually employ a tracking-by-detection method, and data association plays a critical role. This paper proposes a learning and graph-optimized (LEGO) modular tracker to improve data association performance in the existing literature. The proposed LEGO tracker integrates graph optimization, which efficiently formulate the association score map, facilitating the accurate and efficient matching of objects across time frames. To further enhance the state update process, the Kalman filter is added to ensure consistent tracking by incorporating temporal coherence in the object states. Our proposed method utilizing LiDAR alone has shown exceptional performance compared to other online tracking approaches, including LiDAR-based and LiDAR-camera fusion-based methods. LEGO ranked $3^{rd}$ among all trackers (both online and offline) and $2^{nd}$ among all online trackers in the KITTI MOT benchmark for cars [^1] https://www.cvlibs.net/datasets/kitti/eval_tracking.php, at the time of submitting results to KITTI object tracking evaluation ranking board.
@@ I Introduction
Multi-object tracking (MOT) is a crucial technology utilized in various systems within the field of intelligent transportation systems (ITS), like traffic monitoring [1, 2, 3], advanced driver assistance systems (ADAS) [4] and autonomous driving (AD) [5]. Fig. 1 illustrates the task of LiDAR-based MOT, which presents several challenges that need to be addressed:
  • Refinement of state information: The tracker should refine the state information provided by preceding modules, such as instance segmentation [6][7] and object detection [8][9][10]. By incorporating this refined state information, the tracker contributes to the overall accuracy and efficacy of the tracking system.
  • False detection rejection: To enhance the overall reliability of the tracking process, it is essential for the tracker to possess the capability to discern and disregard inaccurate detections produced by the object detector. This ability to filter out false positives is crucial for maintaining accurate tracking results.
  • Identification (ID) estimation and tracking continuity: Another significant challenge in multi-object tracking is maintaining tracking continuity even in situations where the object fails to be detected across successive frames. The tracker should be able to handle occlusions, temporary disappearances, or other situations that can interrupt object detection, ensuring robust and uninterrupted monitoring of the objects. This entails assigning and preserving a unique identification marker for each object, ensuring its continuity and consistency.
@@
Effectively addressing the challenges in MOT is paramount for developing robust and accurate tracking systems. The existing MOT methods can be categorized into distinct groups based on the sensor modalities they utilize. These categories encompass 2D MOT methods, 3D MOT techniques relying solely on LiDAR sensors, and 3D MOT methods that leverage both camera and LiDAR systems. In the realm of 2D MOT, notable methods, such as FairMOT [11], ByteTrack [12], and StrongSort [13] predominantly employ camera data for tracking objects. Although these methods provide valuable insights, their performance is constrained by the limitations of working with 2D representations alone. On the other hand, 3D MOT techniques exclusively employ LiDAR sensors to capture 3D information about the tracked objects. AB3DMOT [20] serves as an exemplary method within this category, demonstrating the efficacy of LiDAR-based approaches for precise object localization and spatial understanding. Furthermore, there exist 3D MOT methods that capitalize on the synergies between camera and LiDAR systems. Probabilistic CBMOT [47] represents one such approach, where both sensor modalities are utilized to improve tracking performance. By combining the strengths of cameras and LiDAR sensors, these methods aim to achieve enhanced object tracking accuracy. A majority of these MOT methods can be classified as tracking-by-detection methods, wherein 3D detectors play a central role. Various 3D detectors have been developed to facilitate object detection in the MOT pipeline. Examples include PointRCNN [14], PointGNN [15], CenterPoint [16], and CasA [17], among others.
(a) Tracking result of time $t-1$
(b) Tracking result of time $t$
Figure 1: Illustration of a MOT method via detection utilizing LiDAR for ADAS and AD: This figure delineates the tracking result at two successive time frames $t-1$ and $t$, shown in birds eye’s view (BEV), the different colours show different tracked objects.
In this research, the tracking-by-detection task for MOT with a specific focus on utilizing LiDAR is adopted. The proposed approach consists of a two-stage process, which involves an initial object detection step followed by a subsequent tracking phase. During the object detection stage, the detector analyzes the input frames and generates relevant information, such as bounding box dimensions and object centre location in 3D. These outputs provide valuable cues about the location and extent of the detected objects within the 3D space. Following the object detection stage, the tracking phase is employed to refine and synthesize the acquired data. In this phase, a tracking algorithm is applied to associate and maintain the IDs of the objects across different frames. Specifically, the tracker assigns unique track IDs to each bounding box, enabling consistent identification of objects throughout the tracking process.
The tracking phase is initiated by generating pairwise affinities, which play a crucial role in associating objects across time frames. Various methods have been employed to compute these affinities in the context of MOT, including feature representation techniques, such as appearance similarity, distance metrics, and motion predictions. In recent years, deep learning-based features rooted in convolutional neural networks (CNN) [43] have gained widespread adoption for computing these affinities. Additionally, a multi-cue approach [64] which incorporates motion as an additional cue, has been adopted to complement appearance features for data association. While motion modules based on deep learning, such as long short-term memory (LSTM) [65], have shown limited robustness and efficiency when compared to motion prediction methods using the Kalman filter [21], relying solely on intersection over union (IoU) and Kalman filter introduces certain challenges. Firstly, there is a loss of point cloud information, and secondly, geometric similarity computed solely based on IoU lacks reliability. Although some studies have employed deep learning-based methods, such as [24], to derive the cost matrix, the noise-corrupted output from the detector limits the feature learning capabilities of deep neural networks. Moreover, the trajectory representation, which consists of a topology unsuited for CNN or recurrent neural networks (RNN), poses challenges when dealing with unstructured data. Considering the trajectory as an edge comprising multiple object nodes suggests that a graph neural network (GNN) can be a suitable deep learning model for this task. However, the direct application of GNN presents certain issues, including the compromised calculation efficiency due to the high dimensionality of the graph and the information redundancy caused by uniform message passing from neighboring nodes [69].
To tackle the aforementioned challenges, an appealing approach is by redefining the data association problem as a bipartite graph matching problem, which takes advantage of its graphical formulation. In this redefined context, a self-attention-based GNN combined with graph structure information is integrated with the Kalman filter to facilitate data association and perform state updates.
Our approach incorporates a new GNN optimization module, which operates by reasoning over non-Euclidean graph data through message passing and combining the graph structure information. This module enhances the relational reasoning abilities of the system and reduces information redundancy. By leveraging self-attention mechanisms and graph structure information, our GNN can effectively capture long-range dependencies and complex relationships between objects, contributing to improved data association accuracy. Additionally, an offset module is specifically designed to refine the estimated bounding boxes from the detector. The primary objective of this module is to correct the output of the detector, resulting in more accurate and reliable information about the object’s motion.
@@
  • A learning and graph-optimized (LEGO) modular tracker is proposed for online MOT. Compared to previous model-based trackers, LEGO introduces two learning modules as its core, an offset correction module and a new GNN, which are easily comprehensible, tunable, and adaptable to different sensor modalities.
  • The offset correction module is designed to predict the offset between these detection results and the ground truth. The primary purpose of this mechanism is to rectify prediction outcomes, thereby improving the accuracy of the model’s predictive capabilities. The adjency matrix guide graph network (AMGN) is introduced to use the information of the graph structure to do the message passing and updating with self-attention, and directly generate the association score matrix.
  • Empirical evaluation and performance analysis: The paper empirically evaluates the proposed AMGN module, demonstrating its effectiveness in enhancing tracking performance. The proposed LEGO is evaluated using the KITTI MOT benchmark, and the results showcase commendable performance in terms of tracking accuracy and reliability.
@@ II-A2 Deep-learning based Methods
SimTrack [24] and CenterTube [40] introduce end-to-end trainable models for joint detection and tracking, leveraging raw point cloud as input. OGR3MOT [25], Batch3DMOT [26], PolarMOT [41], Rethinking3DMOT propose sophisticated graph structures based on neural message passing, enabling online execution of detection and tracking processes. The ENBP-Tracker [32] integrates GNN into its design, distinguishing itself by combining the network with a belief propagation tracker. This hybrid tracking architecture enhances the robustness and efficacy of the tracking process, showcasing the potential of merging traditional tracking methods with advanced neural network architectures. Intertrack [34] TransMOT [30] adopt transformer structures to generate discriminative object representations for data association. Minkowski-Tracker [35], PC-TCNN [45], ShaSTA [36] employ proposal networks to extract features from various feature maps, facilitating the learning of affinity matrix for point cloud-based MOT.
@@ III The Proposed LEGO Modular Tracker
Within this section, we will introduce the LEGO modular tracker. Firstly, we will present its overall framework, followed by a detailed elaboration of each module. Fig. 2 shows the proposed LEGO tracker, which comprises several modules: motion prediction, offset correction, feature extraction, AMGN score calculation, association, state update, and track management.
III-A The Framework of LEGO Modular Tracker
Initiating the process, the 3D point cloud is fed into a 3D object detector, which is responsible for identifying potential objects within the scene. To augment the accuracy of the detection results, an offset correction module is proposed. The primary function of the offset correction module is to anticipate the offset between the detection result and the ground truth, thereby smoothing the detector’s outputs. Meanwhile, the classical motion prediction module takes the tracked state from the previous timestep and predicts the state of each object in the current timestep by Kalman prediction. Then, following the recent model-based offline tracker, i.e. PC3T [44], the geometry and motion cost matrix $\bm{A}$ is calculated by using the detected state after the offset correction module and the predicted state from the motion prediction module.
To extract the robust 3D point cloud features, the outputs of the previous motion prediction and offset correction modules along with the 3D point cloud are fed into the feature extraction module, which utilizes the PointNet architecture [67] to produce the feature maps. Sequentially, a bipartite graph $\bm{G}=\langle\bm{N},\bm{E}\rangle$ is constructed with the predicted feature map and the detected feature map. Here, ${\bm{N}=(\bm{N}_{det},\bm{N}_{pre})}$ represents the set of nodes for detected objects and predicted objects and edges $\bm{E}$ represent the potential association between detected and predicted objects.
Upon inputting the bipartite graph $\bm{G}$ and the learnable adjacency matrix into the proposed AMGN module, the AMGN score matrix $\bm{B}$ is produced as an output. The AMGN module employs a learnable adjacency matrix, which encapsulates the structural information of the graph. This design enables the model to progressively refine its comprehension of the graph’s architecture. Furthermore, it facilitates the quantification and learning of the relational strengths between any given pair of nodes. The final association cost matrix is then obtained via the proposed method, encapsulating both the AMGN score matrix and the geometry and motion cost matrix. The data association module is utilized to discern the matched pairs. Finally, same as all classical model-based trackers, the association map is fed into the state update module, prompting an update in the object states using Kalman filter, within the output being conveyed to the track management module.
@@ III-B Motion Prediction Module
Recently, LSTM networks have become increasingly popular in MOT methods for predicting changes in the object state [66]. These networks are favoured for their ability to handle sequential data exceptionally well. However, it is important to note that LSTM networks require more time than traditional motion prediction techniques, like those based on the Kalman filter. Certain MOT techniques prefer using a constant velocity motion model rather than LSTM networks. Yet, this model assumes that the object moves at a nearly constant speed, which may not hold true in several real-world situations. As a result, if there are consecutive missed detections, it could lead to considerable errors in motion prediction. To overcome this problem, the constant acceleration (CA) motion model [21] is used, which offers a more precise representation of the object state. In the CA motion model, the object state $\bm{X}_{t-1}$ at a specific temporal instance $t-1$ is represented as $\left[\bm{pos}_{t-1},\bm{v}_{t-1},\bm{\alpha}_{t-1}\right]^{T}$, where $\bm{pos}_{t-1}$ denotes the position, $\bm{v}_{t-1}$ represents the velocity, and $\bm{\alpha}_{t-1}$ indicates the orientation of the object. The prediction of the mean state $\bm{X}_{t|t-1}$ and covariance $\bm{P}_{t|t-1}$ using Kalman prediction equations with the CA motion model is as follows:
$\bm{X}_{t|t-1}=\bm{SA}\bm{X}_{t-1},$ (1)
$\bm{P}_{t|t-1}=\bm{SA}\bm{P}_{t-1}\bm{SA}^{T}+\bm{Q},$ (2)
where $\mathbf{SA}$ is the state transition matrix, $\bm{Q}$ represents the motion noise covariance matrix, $\bm{I}$ denotes the identity matrix, $\bm{O}$ represents the zero matrix, $n$ denotes the state dimension, and $\delta$ and $\mathbf{a}$ are sensor-related hyper-parameters.
Figure 3: The structure of the offset correction module. The detection result, which comprises parameters $x_{p}^{d},y_{p}^{d},z_{p}^{d},w^{d},h^{d},l^{d}$, is inputted into the offset prediction head. The offset prediction head can be construed as a simple neural network, which computes the offset between the ground truth and the obtained detection result.
@@ III-C Offset Correction Module
Our investigation has revealed a significant issue related to the accuracy of the state inferred by the detector, especially when errors arise from the 3D object detector. In order to mitigate this potential source of inaccuracies, the integration of an offset correction module is proposed. This module serves the crucial purpose of rectifying the detection outcomes generated by the 3D object detector, thereby enhancing the overall accuracy of the system. The architecture and components of this offset correction module are depicted in Fig. 3. This module comprises a 3D convolution layer, batch normalization layer, and multi-layer perception (MLP) with the residual connection. The offset obtained from the head is given by
$O=\log\left(\frac{\bm{gt}-\bm{dr}}{\beta}\right),$ (3)
where ${\bm{gt}=\left(x_{p}^{g},y_{p}^{g},z_{p}^{g},w^{g},h^{g},l^{g}\right)}$ represents the ground truth coordinates $(x_{p}^{g},y_{p}^{g},z_{p}^{g})$, width $w^{g}$, height $h^{g}$, length $l^{g}$; ${\bm{dr}=\left(x_{p}^{d},y_{p}^{d},z_{p}^{d},w^{d},h^{d},l^{d}\right)}$ indicates the object detection extracted from the 3D object detector, and $\beta$ is the scaling factor. Through Eq. (3), the discrepancies of object detection are smoothed.

### v6

@@ Abstract
Online Multi-Object Tracking (MOT) plays a pivotal role in autonomous systems. The state-of-the-art approaches usually employ a tracking-by-detection method, and data association plays a critical role. This paper proposes a learning and graph-optimized (LEGO) modular tracker to improve data association performance in the existing literature. The proposed LEGO tracker integrates graph optimization, which efficiently formulates the association score map, facilitating the accurate and efficient matching of objects across time frames. To further enhance the state update process, the Kalman filter is added to ensure consistent tracking by incorporating temporal coherence in the object states to further enhance the state update process. Our proposed method, utilising LiDAR alone, has shown exceptional performance compared to other online tracking approaches, including LiDAR-based and LiDAR-camera fusion-based methods. LEGO ranked $3^{rd}$ among all trackers (both online and offline) and $2^{nd}$ among all online trackers in the KITTI MOT benchmark for cars [^1] https://www.cvlibs.net/datasets/kitti/eval_tracking.php, at the time of submitting results to KITTI object tracking evaluation ranking board. Moreover, our method also achieves competitive performance on the Waymo open dataset benchmark.
@@ I Introduction
Tracking is a crucial technology utilized in various systems within the field of intelligent transportation systems (ITS), like pedestrian tracking [1, 2], drone monitoring [3, 4], target tracking [5], traffic monitoring [6, 7], Advanced Driver Assistance Systems (ADAS) and Autonomous Driving (AD) [8, 9]. To develop robust and accurate tracking systems effectively, various approaches have been investigated by using different sensor modalities, encompassing 2D multi-ojbect-tracking (MOT) methods using camera images, 3D MOT techniques relying solely on LiDAR, and 3D MOT methods that leverage both camera and LiDAR. In the realm of 2D MOT, notable methods, such as FairMOT [10], ByteTrack [11], and StrongSort [12] predominantly employ camera data for tracking objects. Although these methods provide valuable insights, their performance is constrained by the limitations of working with 2D representations alone. On the other hand, 3D MOT techniques exclusively employ LiDAR to capture 3D information about the tracked objects. AB3DMOT [21] serves as an exemplary method within this category, demonstrating the efficacy of LiDAR-based approaches for precise object localization and spatial understanding. Furthermore, there are 3D MOT methods that exploit the synergies between camera and LiDAR systems. Probabilistic CBMOT [48] represents one such approach, where both sensor modalities are utilized to improve tracking performance. By combining the strengths of cameras and LiDAR sensors, these methods aim to achieve enhanced object tracking accuracy. A majority of these MOT methods can be classified as tracking-by-detection methods, in which various 3D object detectors, e.g., PointRCNN [13], PointGNN [14], CenterPoint [15], and CasA [16], etc, have been employed to provide estimated bounding boxes as measurements input to the MOT pipeline. Some work, such as [3, 4], may use drone monitoring to collect 2D video image information and perform the tracking based on a temporal transformer. As drone video is usually two-dimensional, it lacks the rich depth information available from other sources. However, continuous point cloud data is a form of 3D video for real-world representation. This is quite different from conventional 2D video, as it captures spatial structure and depth information directly in three dimensions, enabling more accurate modeling, understanding, and tracking of dynamic scenes in complex environments.
Despite many efforts have been made in this area, MOT systems still suffer from erroneous measurements generated from object detector and inaccurate data association. Specifically, erroneous measurements from object detectors lead to misaligned or imprecise bounding boxes, which degrade the quality of input data for tracking. Inaccurate data association, on the other hand, results in track identity switches or fragmented tracks, making it difficult to maintain consistent object trajectories. Consequently, these issues significantly hinder the overall performance of MOT systems.
To address these challenges, we propose a novel framework that explicitly tackles both issues through the introduction of two key components: the offset correction module and the Adjacency Matrix guided Graph Network (AMGN) score calculation module. The offset correction module is designed to correct the erroneous bounding boxes generated by the detector, resulting in more accurate and reliable object localization. Meanwhile, the AMGN score calculation module redefines the data association problem as a bipartite graph matching problem, leveraging a combination of graph structure information, an LSTM, and a new self-attention structure whose coefficients is derived from the learnable adjacency matrix to enhance relational reasoning and reduce information redundancy. This integration allows the Graph Neural Network (GNN) to effectively capture complex dependencies between measurements and predicted tracks, improving the accuracy of data association. Lastly, by integrating with Kalman filter, our approach facilitates more robust state updates and improves overall tracking performance.
@@
  • A easily comprehensible and tunable LEarning and Graph-Optimized (LEGO) modular tracker is proposed for online MOT. LEGO introduces two learning modules as its core, an offset correction module and a Adjacency Matrix guided Graph Network (AMGN) score calculation module.
  • The offset correction module is designed to predict the offset between these detection results and the ground truth. The primary purpose of this mechanism is to rectify prediction outcomes, thereby improving the accuracy of the model’s predictive capabilities. The AMGN score calculation module which contains LSTM and $K$ AMGN blocks, leverages the LSTM and graph structure information to perform message passing and update node features using the self-attention coefficients derived from a learnable adjacency matrix. By doing so, the AMGN score calculation module can directly computes the association score.
  • Empirical evaluation and performance analysis of the proposed offset correction and AMGN score calculation modules, demonstrate their effectiveness in improving tracking performance, as well as the proposed LEGO tracker’s state-of-the-art performance in both the KITTI MOT benchmark and Waymo 3D MOT benchmark.
@@ II-A2 Deep-learning based Methods
SimTrack [25] and CenterTube [41] introduce end-to-end trainable models for joint detection and tracking, leveraging raw point cloud as input. OGR3MOT [26], Batch3DMOT [27], PolarMOT [42], Rethinking3DMOT propose sophisticated graph structures based on neural message passing, enabling online execution of detection and tracking processes. The ENBP-Tracker [33] integrates GNN into its design, distinguishing itself by combining the network with a belief propagation tracker. This hybrid tracking architecture enhances the robustness and efficacy of the tracking process, showcasing the potential of merging traditional tracking methods with advanced neural network architectures. Intertrack [35] TransMOT [31] adopt transformer structures to generate discriminative object representations for data association. Minkowski-Tracker [36], PC-TCNN [46] and ShaSTA [37], employ proposal networks to extract features from various feature maps, facilitating the learning of affinity matrix for point cloud-based MOT.
@@ II-B MOT with LiDAR and Camera Fusion
II-C MOT with Graph Neural Network based Methods
In recent years, GNN-based tracking methods have attracted considerable research interest, primarily due to their capability to model Multi-Object Tracking (MOT) as a graph structure, allowing object nodes to effectively exchange and aggregate features. GNN3DMOT [52] was among the first to propose a GNN-based 3D MOT framework, integrating both 2D and 3D feature learning. Their method employs a feature extractor to obtain motion and appearance features from both 2D images and 3D point clouds, subsequently feeding these features into a GNN to facilitate feature interaction among objects. Furthermore, [51] introduced an MOT approach comprising two separate graph networks: an appearance graph network and a motion graph network. These two GNNs independently compute similarity measures between detected objects and existing trackers through a four-step updating module. The final affinity matrix is then derived by combining these two similarity scores through a weighted fusion mechanism. Additionally, [53] proposed an iterative Graph Convolutional Network (GCN) clustering method designed to reduce computational complexity. This method ranks generated proposals based on estimated quality scores, effectively maintaining proposal quality while significantly enhancing computational efficiency. Similarly, [54] presented a GCN-based MOT approach, modeling the relationships between existing tracklets and intra-frame detections as a general undirected graph to improve tracking robustness and accuracy. Most existing methods perform message passing and updating with Graph Attention Networks (GAT) to ensure that different nodes contribute differently during the message passing process. However, GAT handles message passing and node updating on a node-by-node basis, which overlooks the global structure of the graph. This limitation may result in the loss of critical information embedded within the overall graph structure. To address this issue, we propose the AMGN block, which incorporates a learnable adjacency matrix into the self-attention mechanism during the message passing and updating process. This design enables the model to capture both local and global structural information, leading to more effective graph representation.
@@ III The Proposed LEGO Modular Tracker
Within this section, we will introduce the LEGO modular tracker. Firstly, we will present its overall framework, followed by a detailed elaboration of each module. Fig. 1 shows the proposed LEGO tracker, which comprises several modules: motion prediction, proposed offset correction, feature extraction, proposed AMGN score calculation, association, state update, and track management.
III-A The Overview Framework of LEGO Modular Tracker
Initiating the process, the 3D point cloud is fed into a 3D object detector. To refine these detection results, an offset correction module is introduced. This module predicts the offset between the detection outputs and the ground truth, smoothing the detector’s predictions. Simultaneously, a motion prediction module uses the tracked state from the previous time step to predict the current state of each object, by using Constant Acceleration (CA) motion model [22].
Next, the outputs from the motion prediction and offset correction modules, along with the 3D point cloud, are passed through a feature extraction module. This module, based on the PointNet architecture [73], generates robust 3D feature maps. Using these feature maps, a bipartite graph $\bm{G}=\langle\bm{N},\bm{E}\rangle$ is constructed, where $\bm{N}$ represents nodes for detected and predicted objects, and $\bm{E}$ represents potential associations between them. The bipartite graph $\bm{G}$ and a learnable adjacency matrix are then input into the AMGN block under the proposed AMGN score calculation module, which outputs the AMGN score matrix $\bm{B}$. Using LSTM and $K$ AMGN blocks with the learnable adjacency matrix, the AMGN score calculation module captures and refines the structural relationships within the graph, quantifying the relational strengths between nodes. The final association cost matrix is derived by combining the AMGN score matrix $\bm{B}$, with the geometry and motion cost matrix $\bm{A}$ which is calculated by following PC3T [45].
The data association module then matches detected and predicted objects based on the association cost matrix. Finally, the resulting association map is passed to the state update module, which updates object states using the standard Kalman filter update. The track management module handles unmatched detected states and predicted states by using classical $M/N$ logic. Specifically, if a predicted state remains unmatched for more than $N_{t}$ consecutive frames, it is removed as leaving the scene. Otherwise, unmatched predicted states are retained, assuming the object may reappear in subsequent frames. The track management module also consider all unmatched objects as potential new objects entering the scene. In order to avoid creating false positive trajectories, new trajectory will not be created for the unmatched detection until it has been continually matched in the next $M_{t}$ frames.
@@ III-B Motion Prediction Module
Certain MOT techniques prefer to use a constant-velocity motion model based on the Kalman Filter. This model assumes that the object moves at a nearly constant speed, which may not hold true in several real-world situations. As a result, consecutive missed detections can lead to considerable errors in motion prediction. To overcome this problem, the constant acceleration (CA) motion model [22] is used based on the Kalman Filter, which offers a more precise representation of the object state. In the CA motion model, the object state $\bm{X}_{t-1}$ at a specific temporal instance t1t-1 is represented as $\left[\bm{pos}_{t-1},\bm{v}_{t-1},\bm{\alpha}_{t-1}\right]^{T}$, where $\bm{pos}_{t-1}$ denotes the position, $\bm{v}_{t-1}$ represents the velocity, and $\bm{\alpha}_{t-1}$ indicates the orientation of the object. The prediction of the mean state $\bm{X}_{t|t-1}$ and covariance $\bm{P}_{t|t-1}$ using Kalman prediction equations with the CA motion model is as follows:
$\bm{X}_{t|t-1}=\bm{S}_{A}\bm{X}_{t-1},$ (1)
$\bm{P}_{t|t-1}=\bm{S}_{A}\bm{P}_{t-1}\bm{S}_{A}^{T}+\bm{Q},$ (2)
where $\bm{S}_{A}$ is the state transition matrix, $\bm{Q}$ represents the motion noise covariance matrix, $\bm{I}$ denotes the identity matrix, $\bm{O}$ represents the zero matrix, $n$ denotes the state dimension, and $\delta$ and $\mathbf{a}$ are sensor-related hyper-parameters.
Figure 2: The structure of the offset correction module. The detection result, which comprises the parameters $x_{p}^{d},y_{p}^{d},z_{p}^{d},w^{d},h^{d},l^{d}$, is input into the offset prediction head. The offset prediction head can be constructed” as a simple neural network, which computes the offset between the ground truth and the obtained detection result.
@@ III-C Offset Correction Module
Our investigation has revealed a significant issue related to the accuracy of the state inferred by the detector, especially when errors arise from the 3D object detector. In order to mitigate this potential source of inaccuracies, an offset correction module is proposed. This module serves the crucial purpose of rectifying the detection outcomes generated by the 3D object detector, thereby enhancing the overall accuracy of the system. The architecture and components of this offset correction module are depicted in Fig. 2. This module comprises a 3D convolution layer, batch normalization layer, and Multi-Layer Perception (MLP) with the residual connection. The offset obtained from the head is $O_{pred}$, and the ground truth offset is:
$O_{gt}=\log\left(\frac{gt-dr}{\beta}\right),$ (3)
where $gt=\left(x_{p}^{g},y_{p}^{g},z_{p}^{g},w^{g},h^{g},l^{g}\right)$} represents the ground truth bounding box coordinates $(x_{p}^{g},y_{p}^{g},z_{p}^{g})$, width $w^{g}$, height $h^{g}$, length $l^{g}$; $dr=\left(x_{p}^{d},y_{p}^{d},z_{p}^{d},w^{d},h^{d},l^{d}\right)$} indicates the object detection bounding box extracted from the 3D object detector, and $\beta$ is the scaling factor.
The offset correction module is directly supervised by the ground-truth offset. The Hungarian algorithm is used to match the ground truth bounding box and the predicted bounding box. Specifically, the loss function is defined for training the offset correction module as the Smooth L1 loss between the predicted offset $O_{pred}$ and the ground-truth offset $O_{gt}$, formulated as follows:
$L_{offset}=\text{SMOOTHL1}\left(O_{pred},O_{gt}\right).$ (4)
The ground truth data is utilized exclusively in the training phase of the offset correction module. Specifically, the module is trained to predict the offset between the ground truth annotations and the detector outputs, allowing it to learn systematic error patterns inherent in the detector. However, at inference time, the ground truth data is not available. Instead, the trained offset correction module directly predicts and rectifies the offsets based solely on the detection results provided by the 3D object detector.